OpenStudy (anonymous):

A collection of quarters and nickels is worth $14.80. There are 88 coins in all. How many of each coin are there? quarters__ nickels__ @StudyGurl14

OpenStudy (studygurl14):

Q = quarters N = nickels 25Q + 5N = 1480 Q + N = 88 Solve using those equations

OpenStudy (studygurl14):

(quarters are worth 25 cents, nickels are 5 cents0

OpenStudy (anonymous):

so 25+5 until i get 1480?@StudyGurl14

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

um...no...Use the system of equations and solve for Q and N

OpenStudy (studygurl14):

25Q + 5N = 1480 Q + N = 88 Q = 88 - N 25(88-N)+5N = 1480 2200 - 25N + 5N = 1480 -20N = -720 N = 36 So..36 nickels Can you do the rest?

OpenStudy (anonymous):

yes i think 52 quarters ?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

Right. :)

OpenStudy (anonymous):

thanks :)

OpenStudy (studygurl14):

anytime :)